LOG-IN
Displaying report 1-1 of 1.
VIS (SR3)
Print this report.
ryohei.kozu - 18:11, Monday 03 December 2018 (7268)Get code to link to this report
Comment to SR3-RM was sitting on the EQ frame (Click here to view original report: 7236)

Several filter damped the SR3. Here is the remeasured TF with BW = 0.01 Hz.

TF of IPs (L, T, Y),

 

of GAS filters (BF, F1, F0),

 

 

of IM (L,T, V, R, P, V).

 

Images attached to this comment
7268_20181203095054_sr3bf.png 7268_20181203095126_sr3f0.png 7268_20181203095152_sr3f1.png 7268_20181203095200_sr3iml.png 7268_20181203095207_sr3imp.png 7268_20181203095314_sr3imr.png 7268_20181203095327_sr3imt.png 7268_20181203095348_sr3imv.png 7268_20181203095358_sr3imy.png 7268_20181203095404_sr3ipl.png 7268_20181203095410_sr3ipt.png 7268_20181203095423_sr3ipy.png